OpenStudy (anonymous):

For which p∈R is the infinite series ∑k=2∞=k(−1)k(k−1)2logp(k) convergent? conditionally convergent? absolutely convergent?

OpenStudy (kc_kennylau):

mind drawing it out?

OpenStudy (kc_kennylau):

Calculate the ratio between the \(n\)th term and the \((n+1)\)th term.

OpenStudy (kc_kennylau):

If \(\left|\dfrac{n+1}n\right|>1\), divergent. If it's <1, convergent.

OpenStudy (kc_kennylau):

if it's =1, conditionally convergent.

OpenStudy (anonymous):

sorry.. it should have been \[\sum_{k=2}^{\infty} =\frac{ k(-1)^{k} }{ (k-1)^{2}\log ^{p} (k)}\] so, kc_kennylau, your comment is still valid ?

OpenStudy (kc_kennylau):

um...

OpenStudy (kc_kennylau):

There are many convergence tests, I only know one of them...

OpenStudy (anonymous):

The answer depends on p.

OpenStudy (anonymous):

For this one use the integral test and the problem that I solved for you before.

OpenStudy (perl):

no its an alternating series, so use alternating series test

OpenStudy (anonymous):

The question is not only about convergence only. Please read the question.

OpenStudy (anonymous):

Try to prove that if p>1, then the series is absolutely convergent. If \( p \le 1\) one can show that is not absolutely convergent but convergent. Hence it is conditionally convergent.

OpenStudy (anonymous):

If the ratio is 1 then it is inconclusive, not convergent.

OpenStudy (anonymous):

Look at the series: \[\sum_n \frac{1}{n}\] The harmonic series is divergent but the ratio test gives 1.

OpenStudy (anonymous):

Also, that is sloppy notation. You are not taking: \[{\lim_{n \rightarrow \infty} \left| \frac{n+1}{n}\right|}\] If your sum is: \[\sum_n a_n\] You're considering: \[\lim_{n \rightarrow \infty} \left| \frac{a_{n+1}}{a_n}\right|<1\] which tells your series is decreasing and therefore convergent.

OpenStudy (anonymous):

So evaluate that limit for your series. From there impose the restriction and you'll find your condition on p. That answers your question.